OpenStudy (wwe123):

the sum of three consecutive terms of ap is 9 and sum of their product 35

OpenStudy (wwe123):

SORRY the sum of three consecutive terms of ap is 9 and sum of their squares is 35 find \[S _{n}\]

OpenStudy (anonymous):

Can you write the three terms of the ap?

OpenStudy (wwe123):

a-d,a,a+d

OpenStudy (anonymous):

Let the 3 terms be (a-d),(a) and (a+d) Here, a-d+a+a+d = 3a = 9 , or a = 3 .. Similarly, do for their squares, and then, do tell what do you get!

OpenStudy (wwe123):

d=2

OpenStudy (wwe123):

then Sn is ..

OpenStudy (hba):

@wwe123 Do you understand anything ?

OpenStudy (wwe123):

i understand well and also able to solve the question,but Answer not match

OpenStudy (hba):

Can you show me your working so i can tell you where you went wrong ?

OpenStudy (wwe123):

1 min

OpenStudy (wwe123):

a-d+a+a+d= 9 3a=9 a=3

OpenStudy (hba):

please continue ?

OpenStudy (anonymous):

You know the Sn formula, no ? n/2(2a+(n-1)d)

OpenStudy (wwe123):

(a-d)^2+a^2+(a-d)^2=35 a^2+d^2-2ad+a^2+a^2+d^2+2ad=35 3a^2+2d^2=35 2d^2=35-27 2d^2=8 d^2=4 d=2

OpenStudy (anonymous):

Now you know "a" and "d" , put these in the Sn formula .. Don't you get the answer that way ?? :/

OpenStudy (wwe123):

i get S=n(2+n)

OpenStudy (wwe123):

@hba

OpenStudy (wwe123):

but answer is n^2 or n(6-n)

hartnn (hartnn):

well, the terms are 1,3,5 for which neither n^2 nor n(6-n) is correct

hartnn (hartnn):

S= n/2(2*(3-2)+(n-1)*2) = n(1+n-1) = n(n) = n^2

OpenStudy (wwe123):

i think given anwer is wrong becoz Sn=n^2 when a=2,d=2

hartnn (hartnn):

the first term is a-d = 3-2 = 1 then u get n^2.

OpenStudy (wwe123):

yes

hartnn (hartnn):

with d^2 = 4, u also get d=-2

hartnn (hartnn):

S= n/2(2*(3+2)+(n-1)*(-2)) = n(5-n+1) = n(6-n) there u go....

hartnn (hartnn):

here d was -2 so , first term = a-d = 3-(-2) = 3+2